# Minimax-Regret Sample Selection in Randomized Experiments

### v1

@@ Abstract
Randomized controlled trials (RCTs) are often run in settings with many subpopulations that may have differential benefits from the treatment being evaluated. We consider the problem of sample selection, i.e., whom to enroll in an RCT, such as to optimize welfare in a heterogeneous population. We formalize this problem within the minimax-regret framework, and derive optimal sample-selection schemes under a variety of conditions. We also highlight how different objectives and decisions can lead to notably different guidance regarding optimal sample allocation through a synthetic experiment leveraging historical COVID-19 trial data.
@@ 1 Introduction
Perhaps our most surprising finding is as follows. First let $U_{g}$ denote the $g$-th group’s expected utility. Suppose we can assign different treatments to different groups, and seek to optimize the weighted average utility $U=\sum_{g=1}^{G}\alpha_{g}U_{g}$ where the weights $\alpha_{g}$ are proportional to the size of each group in the population. [^1] This aggregated utility objective is a scaled version of utilitarian social welfare that weights everyone in the population equally (Moulin, 2004). Then, the minimax-regret study design allocates samples to the $g$-th group proportionally to $\alpha_{g}^{2/3}N$. In particular, this recommendation deviates from the conventional practice of selecting samples in a stratified experiment in proportion to each group’s share in the population (Singh and Masuku, 2013; Sharma, 2017), and instead over-samples minority groups. In Section 3, we discuss reasons for this finding. We also prove that more familiar sample-allocation schemes, like proportional sampling, are minimax optimal under alternate decision theoretic framings. We also then do a synthetic simulation using historical COVID-19 trial data, which illustrates the notably different sample allocations that arise under different framings (Section 4). Overall, our results highlight how different welfare-focused objectives have different implications as to optimal sample selection.
@@ 1.1 Related work
Our research question falls within literature on using decision theory to guide the design of RCTs. As noted above, Manski and Tetenov (2016, 2019) and Azevedo et al. (2020, 2023) use decision theoretic criteria to choose the sample size in an RCT. Another line of work compares different randomization rules over the treatment and control sample allocation in terms of the resulting variance for the average treatment effect, and seeks intervention randomization rules with minimax variance (Bai, 2021; Kallus, 2021; Li, 1983; Wu, 1981). Banerjee et al. (2020) study the value of randomization itself under a number of different decision theoretic frameworks. However, we are not aware of prior work in this literature that considers the question of sample allocation across groups.
Finally, our work also relates to the broader literature on fairness in data analysis. RCTs have often been run in blatantly unfair ways. For example, until recently, most medical research in the United States was conducted on white men, while excluding women and racial minorities (Dresser, 1992); and FDA-approved trials still under-sample black participants relative to their share of the population (Alsan et al., 2022). While it is easy to agree that some ways of running randomized trials are unfair, it can be much harder to say what makes a trial fair. For example, if there is a minority subgroup that has the potential for both disproportionate benefits and disproportionate risks from participating in a study, what considerations should go into deciding how to fairly include them in the study? In such settings results on welfare-driven sample selection may help provide useful insights into potential desiderata for fair RCT design.
@@ 2 Minimax-Regret Sample Selection
Below, we give a near-optimal solution to this minimax-regret problem when the data-generating distribution $D$ is from a Gaussian [^2] When the data-generating distribution is generalized to be from a sub-Gaussian class, a an asymptotic limits-of-experiments analogue to Theorem 2.1 holds in the large sample limit; see Hirano and Porter (2009) for a general discussion of results of this type. class. The sub-optimality in our result is only due to rounding to achieve integer sample allocations, and our solution is optimal whenever no rounding is needed in (5). We present a brief proof of the theorem, with proof of the technical lemmas deferred to the supplementary material. Compared with a conventional proportional selection, the minimax selection rule dilutes the influence of the weights $\alpha_{g}$, and oversamples the minority groups.
@@
Suppose $Y_{g,i}(0)$ and $Y_{g,i}(1)$ are generated i.i.d. from Gaussian distributions $N(b_{g}-\tau_{g}/2,s_{0,g}^{2})$ and $N(b_{g}+\tau_{g}/2,s_{1,g}^{2})$, respectively. Suppose that $s_{0,g}^{2},s_{1,g}^{2},\alpha_{g}$ are bounded away from 00 and $\infty$ for all $g$. Then among 1:1 completely stratified designs with at most $N$ total units, the sample selection $\mathbf{n}^{*}=\left(n^{*}_{1},n^{*}_{2},\cdots,n^{*}_{G}\right)^{\top}$ with
@@
[…]
@@
We start by verifying the last part of the statement, i.e., that given the available data we should choose a decision based on the decision rule induced by a threshold of the difference-in-means estimator. This result is well known; see Stoye (2012) and Tetenov (2012) for a discussion.
@@
[…]
@@
Our proposed sample allocation rule $n^{*}_{g}$ is obtained by rounding $\tilde{n}_{g}$. Because $\mathcal{N}\subset\widetilde{\mathcal{N}}$, we have
$C_{0}\sum_{g=1}^{G}\alpha_{g}\sqrt{\frac{2(s_{0,g}^{2}+s_{1,g}^{2})}{\tilde{n}_{g}}}\leq\min_{\mathbf{n}\in{\mathcal{N}}}\inf_{\delta(D)}\max_{\tau\in\mathbb{R}^{G}}\mathbb{E}_{D}\left[R(\mathbf{n},\delta(D))\right].$ (11)
Notice that $\tilde{n}_{g}-2<n^{*}_{g}\leq\tilde{n}_{g}$. As a result,
$\begin{split}&\inf_{\delta(D)}\max_{\tau\in\mathbb{R}^{G}}\mathbb{E}_{D}\left[R(\mathbf{n}^{*},\delta(D))\right]-\min_{\mathbf{n}\in\mathcal{N}}\inf_{\delta(D)}\max_{\tau\in\mathbb{R}^{G}}\mathbb{E}_{D}\left[R(\mathbf{n},\delta(D))\right]\\ &\qquad\qquad\leq C_{0}\sum_{g=1}^{G}\alpha_{g}\sqrt{\frac{2(s_{0,g}^{2}+s_{1,g}^{2})}{n^{*}_{g}}}-C_{0}\sum_{g=1}^{G}\alpha_{g}\sqrt{\frac{2(s_{0,g}^{2}+s_{1,g}^{2})}{\tilde{n}_{g}}}\\ &\qquad\qquad=C_{0}\sum_{g=1}^{G}\alpha_{g}\sqrt{2(s_{0,g}^{2}+s_{1,g}^{2})}\left((n^{*}_{g})^{-1/2}-(\tilde{n}_{g})^{-1/2}\right)\\ &\qquad\qquad=C_{0}\sum_{g=1}^{G}\alpha_{g}\sqrt{2(s_{0,g}^{2}+s_{1,g}^{2})}\left((n^{*}_{g})^{-1/2}-(n^{*}_{g})^{-1/2}(1+\frac{\tilde{n}_{g}-n^{*}_{g}}{n^{*}_{g}})^{-1/2}\right)\\ &\qquad\qquad=C_{0}\sum_{g=1}^{G}\alpha_{g}\sqrt{2(s_{0,g}^{2}+s_{1,g}^{2})}\left((n^{*}_{g})^{-1/2}-(n^{*}_{g})^{-1/2}\sum_{j=0}^{\infty}\binom{-1/2}{j}\left(\frac{\tilde{n}_{g}-n^{*}_{g}}{n^{*}_{g}}\right)^{j}\right)\\ &\qquad\qquad\leq C_{0}\sum_{g=1}^{G}\alpha_{g}\sqrt{2(s_{0,g}^{2}+s_{1,g}^{2})}(n^{*}_{g})^{-1/2}\left(\frac{\tilde{n}_{g}-n^{*}_{g}}{2n^{*}_{g}}\right)\\ &\qquad\qquad=\mathcal{O}\left(N^{-3/2}\right).\end{split}$ (12)
where the third equality uses the generalized binomial inequality, and the second inequality upper bounds the sum. From Lemmas 2.2 and 10, $\min_{\mathbf{n}\in\mathcal{N}}H\left(\mathbf{n}\right)=\Theta\left(N^{-1/2}\right)$. Combining this with (12) gives rise to the claimed result. ∎
@@ 3 Conventional Sample Selections as Minimax-Regret Solutions
Above, we found that the minimax-regret sample selection suggests oversampling minority groups and employing a sample allocation proportional to $\alpha_{g}^{2/3}$, which is not a common choice in the literature—at least currently. In this section, we discuss two widely adopted sample selection methods, proportional selection and equal selection. We demonstrate that these common selections can also be regarded as regret-minimax selections when the experimenter makes a single treatment decision for all groups, or, respectively, uses an alternate utility function from that we presented in Section 2.
@@ 3.1 Proportional Selection
$\begin{split}&\tilde{\delta}^{DM}(D)=I(\hat{\tau}^{\text{DM,p}}\geq 0),\\ &\hat{\tau}^{\text{DM,p}}=\frac{2}{N}\sum_{g=1}^{G}\sum_{\left\{i:W_{g,i}=1\right\}}Y_{g,i}-\frac{2}{N}\sum_{g=1}^{G}\sum_{\left\{i:W_{g,i}=0\right\}}Y_{g,i}.\end{split}$ (13)
@@ 3.2 An Egalitarian Solution
One may notice that the resulting sample selection shares a common ground with the Neyman allocation (also known as the optimal allocation), which advocates sampling proportionally to the standard deviation (Neyman, 1934). Indeed, if the objective in Neyman’s optimal allocation problem is instead the variance in the worst-off group rather than the sum of variance over all groups, the optimal selections given by the two problems would be the same.
@@ 4 Case Study: A COVID-19 Vaccine Trial
To further explore the implications of different group sample size allocation practices, we conduct a semi-synthetic analysis using historical data from COVID-19 vaccine development. Different demographic subpopulations may experience different benefits and side effects from vaccines, and so it is of interest to consider experimental design to inform vaccination recommendations for subpopulations. We recognize that for real-world clinical trials there are many important considerations taken into account when designing an experimental trial, and here we use this simply to help illustrate how different decision-theoretic criteria may lead to notably different group allocations.
We use data from a large-scale Phase 3 randomized, placebo-controlled COVID-19 vaccine trial conducted at 99 centers across the United States (Baden et al., 2021). To account for anticipated heterogeneity in the treatment effect across different demographic groups, the randomization is stratified into three groups: 1. $\geq 18$ and $<65$ years and not at risk (for severe Covid-19); 2. $\geq 18$ and $<65$ years and at risk; 3. $\geq 65$ years. In an effort to demonstrate how the proposed sample selection schemes can be used in practice, we explore the sample selections chosen with the three minimax sample selections given in Theorems 2.1, 16 and 19, and compare their performances under both the worst-case incidence rates and the incidence rates reported in Baden et al. (2021).
@@ 4.1 Variance Approximation
According to CDC’s data recorded as of September 2020 (during the period when the protocol was being prepared), the incidence rate of new admissions of patients with confirmed COVID-19 was around $0.7\%$ for people $>=18$ to $<65$ yr, and $2.5\%$ for people $>=65$ yr (Centers for Disease Control and Prevention, 2020). Furthermore, according to a phase 1 trial of an mRNA vaccine, the incidence rate of severe solicited adverse reactions was around $6.7\%$ in a dose group of 100mcg (Jackson et al., 2020). Given this, we consider a conservative approximation of the noise levels, where we take the incidence rates of severe COVID-19 in the treated group to be the same as the ones in the control group and the incidence rates of severe solicited adverse reactions in the control group to be the same as the ones in the treated group. Those approximated noise levels are summarized in Table 1.
@@ 4.2 Budget Constraint
While there is no explicit requirement that sample sizes must be based on power calculations, it is generally considered good practice to conduct power calculations to determine the appropriate sample size for clinical trials (Food and Drug Administration and others, 1988). Thus, we determine the total sample size constraint by adopting the approach outlined in Baden et al. (2021), which sets $N$ to be the sample size required to achieve a power of 0.9 under a specified hypothesis test.
@@
To calculate the total sample size, we assume that $\tilde{Y}_{g,i,1}(1)$ and $\tilde{Y}_{g,i,1}(0)$ are generated i.i.d. from $N(b_{\text{COVID}}-\tau_{\text{COVID}}/2,s_{0,\text{COVID}}^{2})$ and $N(b_{\text{COVID}}+\tau_{\text{COVID}}/2,s_{1,\text{COVID}}^{2})$, respectively, with $s_{w,\text{COVID}}^{2}=\sum_{g=1,2}\alpha_{g}s_{w,g,\text{COVID}}^{2}$, $s_{w,g,\text{COVID}}$ denoting the standard deviations of $Y_{g,i,1}(w)$, and $b_{\text{COVID}}$ being a nuisance baseline parameter. We find the required total sample size $N$ that, with a difference-in-means estimator,
$\mathbb{P}\left[\frac{\hat{\tau}_{\text{COVID}}^{\text{DM}}}{\sqrt{\operatorname{Var}\left[\hat{\tau}_{\text{COVID}}^{\text{DM}}\right]}}\leq Z_{5\%}\right]=90\%,$ (21)
using the approximated variances calculated in Section 4.1. Note that
@@
$N=\frac{2(s_{0,\text{COVID}}^{2}+s_{1,\text{COVID}}^{2})\left(Z_{80\%}-Z_{5\%}\right)^{2}}{{0.60\%}^{2}}\approx 9320.$ (23)

### v2

@@ Abstract
Randomized controlled trials are often run in settings with many subpopulations that may have differential benefits from the treatment being evaluated. We consider the problem of sample selection, i.e., whom to enroll in a randomized trial, such as to optimize welfare in a heterogeneous population. We formalize this problem within the minimax-regret framework, and derive optimal sample-selection schemes under a variety of conditions. Using data from a COVID-19 vaccine trial, we also highlight how different objectives and decision rules can lead to meaningfully different guidance regarding optimal sample allocation.
@@ 1 Introduction
Perhaps our most surprising finding is as follows. First let $U_{g}$ denote the $g$-th group’s expected utility. Suppose we can assign different treatments to different groups, and seek to optimize the weighted average utility $U=\sum_{g=1}^{G}\alpha_{g}U_{g}$ where the weights $\alpha_{g}$ are proportional to the size of each group in the population. [^1] This aggregated utility objective is a scaled version of utilitarian social welfare that weights everyone in the population equally (Moulin, 2004). Then, the minimax-regret study design allocates samples to the $g$-th group proportionally to $\alpha_{g}^{2/3}N$. In particular, this recommendation deviates from the conventional practice of selecting samples in a stratified experiment in proportion to each group’s share in the population (Singh and Masuku, 2013; Sharma, 2017), and instead over-samples minority groups. In Section 3, we discuss reasons for this finding. We also prove that more familiar sample-allocation schemes, like proportional sampling, are minimax optimal under alternate decision theoretic framings. We also consider a hypothetical application of our method guided by historical COVID-19 trial data, which illustrates the notably different sample allocations that arise under different framings (Section 4). Overall, our results highlight how different welfare-focused objectives have different implications as to optimal sample selection.
@@ 1.1 Related work
Our research question falls within literature on using decision theory to guide the design of RCTs. As noted above, Manski and Tetenov (2016, 2019) and Azevedo et al. (2020, 2023) use decision theoretic criteria to choose the sample size in an RCT. Another line of work compares different randomization rules over the treatment and control sample allocation in terms of the resulting variance for the average treatment effect, and seeks intervention randomization rules with minimax variance (Bai, 2023; Kallus, 2021; Li, 1983; Wu, 1981). Banerjee et al. (2020) study the value of randomization itself under a number of different decision theoretic frameworks. However, we are not aware of prior work in this literature that considers the question of sample allocation across groups.
Finally, our work also relates to the broader literature on fairness in data analysis. RCTs have often been run in blatantly unfair ways. For example, until recently, a substantial fraction of medical research in the United States was conducted on white men, while excluding women and racial minorities (Dresser, 1992); and FDA-approved trials still under-sample black participants relative to their share of the population (Alsan et al., 2022). However, while it is easy to agree that some ways of running randomized trials are unfair, it can be much harder to say what makes a trial fair. For example, if there is a minority subgroup that has the potential for both disproportionate benefits and disproportionate risks from participating in a study, what considerations should go into deciding how to fairly include them in the study? In such settings results on welfare-driven sample selection may help provide useful insights into potential desiderata for fair RCT design. Similar considerations are also relevant to fairness-aware data collection when training machine learning and AI systems (Holstein et al., 2019).
@@ 2 Minimax-Regret Sample Selection
Below, we give a near-optimal solution to this minimax-regret problem when the data-generating distribution $D$ is from a Gaussian [^2] When the data-generating distribution is generalized to be from a sub-Gaussian class, an asymptotic limits-of-experiments analogue to Theorem 2.1 holds in the large sample limit; see Hirano and Porter (2009) for a general discussion of results of this type. class. The sub-optimality in our result is only due to rounding to achieve integer sample allocations, and our solution is optimal whenever no rounding is needed in (5). We present a brief proof of the theorem, with proof of the technical lemmas deferred to the supplementary material. Compared with a conventional proportional selection, the minimax selection rule dilutes the influence of the weights $\alpha_{g}$, and oversamples the minority groups.
@@
Suppose $Y_{g,i}(0)$ and $Y_{g,i}(1)$ are generated i.i.d. from Gaussian distributions $N(b_{g}-\tau_{g}/2,s_{0,g}^{2})$ and $N(b_{g}+\tau_{g}/2,s_{1,g}^{2})$, respectively. [^3] For simplicity, we assume the experimenter knows the standard errors. However, $s_{0,g}$ and $s_{1,g}$ can also be regarded as upper bounds on unknown standard errors, as the adversary would maximize the noise levels. Suppose that $s_{0,g}^{2},s_{1,g}^{2},\alpha_{g}$ are bounded away from 00 and $\infty$ for all $g$. Then among 1:1 completely stratified designs with at most $N$ total units, the sample selection $\mathbf{n}^{*}=\left(n^{*}_{1},n^{*}_{2},\cdots,n^{*}_{G}\right)^{\top}$ with
@@
$\begin{split}&\delta^{DM}(D)=\left(I(\hat{\tau}^{\text{DM}}_{1}>0),\cdots,I(\hat{\tau}^{\text{DM}}_{G}>0)\right)^{\top}\\ &\hat{\tau}^{\text{DM}}_{g}=\frac{2}{n_{g}}\sum_{\left\{i:W_{g,i}=1\right\}}Y_{g,i}-\frac{2}{n_{g}}\sum_{\left\{i:W_{g,i}=0\right\}}Y_{g,i},\qquad\qquad g=1,\dots,G\end{split}$ (7)
@@
We start by verifying the last part of the statement, i.e., that given the available data we should choose a decision based on the decision rule induced by a threshold of the difference-in-means estimator, and the induced regret is linear in standard errors. This result is common in statistical treatment choice with Gaussian errors; see Stoye (2012) and Tetenov (2012) for a discussion.
@@
$\begin{split}&\mathbb{E}_{D}\left[R(\mathbf{n},\delta^{\text{DM}}(D))\right]\leq C_{0}\sum_{g=1}^{G}\alpha_{g}\sqrt{\frac{2(s_{0,g}^{2}+s_{1,g}^{2})}{n_{g}}}\end{split}$ (9)
@@
Our proposed sample allocation rule $n^{*}_{g}$ is obtained by rounding $\tilde{n}_{g}$.
Lemma 2.4.
Under the assumptions in Theorem 2.1,
$\begin{split}\inf_{\delta(D)}\max_{\tau\in\mathbb{R}^{G}}\mathbb{E}_{D}\left[R(\mathbf{n}^{*},\delta(D))\right]-\min_{\mathbf{n}\in\mathcal{N}}\inf_{\delta(D)}\max_{\tau\in\mathbb{R}^{G}}\mathbb{E}_{D}\left[R(\mathbf{n},\delta(D))\right]=\mathcal{O}\left(N^{-3/2}\right).\end{split}$ (11)
From Lemmas 2.2 and 10, $\min_{\mathbf{n}\in\mathcal{N}}H\left(\mathbf{n}\right)=\Theta\left(N^{-1/2}\right)$. Combining this with (11) gives rise to the claimed result. ∎
@@ 3 Conventional Sample Selections as Minimax-Regret Solutions
Above, we found that the minimax-regret sample selection suggests oversampling minority groups and employing a sample allocation proportional to $\alpha_{g}^{2/3}$, which is not a common choice in the literature—at least currently. In this section, we discuss two widely adopted sample selection methods, proportional selection and equal selection. We demonstrate that these common selections can also be regarded as regret-minimax selections when the experimenter makes a single treatment decision for all groups, or, respectively, uses an alternate utility function from the one that we used in Section 2.
@@ 3.1 Proportional Selection
$\begin{split}&\tilde{\delta}^{DM}(D)=I(\hat{\tau}^{\text{DM,p}}>0),\\ &\hat{\tau}^{\text{DM,p}}=\frac{2}{N}\sum_{g=1}^{G}\sum_{\left\{i:W_{g,i}=1\right\}}Y_{g,i}-\frac{2}{N}\sum_{g=1}^{G}\sum_{\left\{i:W_{g,i}=0\right\}}Y_{g,i}.\end{split}$ (12)
@@ 3.2 An Egalitarian Solution
One may notice that the resulting sample selection shares a common ground with the Neyman allocation (also known as the optimal allocation), which advocates sampling proportionally to the standard deviation (Neyman, 1934). However, Neyman’s goal was to minimize the variance of a sample mean estimator, whereas our focus is on the quality of the decision and we target directly the regret with egalitarian social welfare.
@@ 4 Case Study: A COVID-19 Vaccine Trial
To further explore the implications of different group sample size allocation practices, we conduct a semi-synthetic analysis using historical data from COVID-19 vaccine development. Different demographic subpopulations may experience different benefits and side effects from vaccines, and so it is of interest to consider experimental design to inform vaccination recommendations for subpopulations. In real-world clinical trials, there are many important considerations beyond the ones discussed in this paper; our goal here is simply to help illustrate how different decision-theoretic criteria may lead to notably different group allocations.
We use data from a large Phase 3 randomized, placebo-controlled COVID-19 vaccine trial conducted at 99 centers across the United States (Baden et al., 2021). To account for anticipated heterogeneity in the treatment effect across different demographic groups, the randomization is stratified into three groups: 1. $\geq 18$ and $<65$ years and not at risk (for severe Covid-19); 2. $\geq 18$ and $<65$ years and at risk; 3. $\geq 65$ years. In an effort to demonstrate how the proposed sample selection schemes can be used in practice, we explore the sample selections chosen with the three minimax sample selections given in Theorems 2.1, 15 and 18, and compare their performances under both the worst-case incidence rates and the incidence rates reported in Baden et al. (2021).
@@ 4.1 Variance Approximation
According to CDC’s data recorded as of September 2020 (during the period when the protocol was being prepared), the incidence rate for newly admitted patients with confirmed COVID-19 was around $0.7\%$ for people $>=18$ to $<65$ yr, and $2.5\%$ for people $>=65$ yr (Centers for Disease Control and Prevention, 2020). Furthermore, according to a phase 1 trial of an mRNA vaccine, the incidence rate of severe solicited adverse reactions was around $6.7\%$ in a dose group of 100mcg (Jackson et al., 2020). Given this, we consider a conservative approximation of the noise levels, where we take the incidence rates of severe COVID-19 in the treated group to be the same as the ones in the control group and the incidence rates of severe solicited adverse reactions in the control group to be the same as the ones in the treated group. Those approximated noise levels are summarized in Table 1.
@@ 4.2 Budget Constraint
While there is no explicit requirement that sample sizes must be based on power calculations, it is generally considered good practice to conduct power calculations to determine the appropriate sample size for clinical trials (Food and Drug Administration, 1988). Thus, we determine the total sample size constraint by following Baden et al. (2021) that sets $N$ to be the sample size required to achieve a power of 0.9 under a specified one-sided hypothesis test with $\alpha=0.05$.
@@
To calculate the total sample size, we adopt the approach outlined in Charan and Biswas (2013) and assume that $\tilde{Y}_{g,i,1}(1)$ and $\tilde{Y}_{g,i,1}(0)$ are generated i.i.d. from $N(b_{\text{COVID}}-\tau_{\text{COVID}}/2,s_{0,\text{COVID}}^{2})$ and $N(b_{\text{COVID}}+\tau_{\text{COVID}}/2,s_{1,\text{COVID}}^{2})$, respectively, with $s_{w,\text{COVID}}^{2}=\sum_{g=1,2}\alpha_{g}s_{w,g,\text{COVID}}^{2}$, $s_{w,g,\text{COVID}}$ denoting the standard deviations of $Y_{g,i,1}(w)$, and $b_{\text{COVID}}$ being a nuisance baseline parameter. We find the required total sample size $N$ that, with a difference-in-means estimator,
$\mathbb{P}\left[\frac{\hat{\tau}_{\text{COVID}}^{\text{DM}}}{\sqrt{\operatorname{Var}\left[\hat{\tau}_{\text{COVID}}^{\text{DM}}\right]}}>Z_{95\%}\,\big{|}\,H_{1}\right]=90\%.$ (20)
Note that under $H_{1}$,
@@
$N=\frac{2(s_{0,\text{COVID}}^{2}+s_{1,\text{COVID}}^{2})\left(Z_{90\%}+Z_{95\%}\right)^{2}}{{\mathbb{E}\left[\hat{\tau}^{\text{DM}}_{\text{COVID}}\,\big{|}\,H_{1}\right]}^{2}}\approx 9320.$ (22)